An object takes 29.5 years to orbit the Sun. What is its average distance from the Sun?

a. 29.5 AU

b. 14.75 AU

c. 9.55 AU
Answer #1

using keplers 3rd law of planetary motion

T^2 = 4 pi^2 R^3 / ( GM)

(29.5* 365* 24*3600)^2 = 4*3.14^2 * R^3 / (6.67*10^-11* 1.99*10^30)

R = 9.55 AU

otpion (c) is correct

===========

do comment in case any doubt, will reply for sure.. Goodluck